OpenStudy (hannahc234):

CHECK MY ANSWER PLEASE!!! WILL FAN AND MEDAL!!!! What is the solution of n^2-49=0 A. -7 B. 7 C. ±7 <----- D. no solution What is the solution of X^2+64=0 A. -5 B. 8 C. ±8 D. No solution <----- What is the side length of a square with the area of 144x^2? A. 12 B. 12x <------ C. ±12x D. No solution

OpenStudy (anonymous):

all are correct.

OpenStudy (janu16):

number 1 is correct number 2 is wrong number is correct

OpenStudy (janu16):

@surjithayer isnt number 2 + or - 8

OpenStudy (anonymous):

ya he is right

OpenStudy (janu16):

@HannahC234 according to me i think number 2 is c

OpenStudy (hannahc234):

I thought you couldn't find a square root for a negative number?

OpenStudy (janu16):

who?@lovefootball

OpenStudy (anonymous):

\[x^2=-64=\left( 8 \iota \right)^2,x=\pm 8 \iota\]

OpenStudy (anonymous):

you

OpenStudy (janu16):

@HannahC234 evryone agrees with C

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

\[ 8 \iota \] is not real so no real solution.

OpenStudy (hannahc234):

So I was right?

OpenStudy (anonymous):

no

OpenStudy (janu16):

he is saying its not real so idk

OpenStudy (anonymous):

it is c

OpenStudy (hannahc234):

Okay thanks

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

It is D.

OpenStudy (hannahc234):

Are you sure?

OpenStudy (hannahc234):

@surjithayer you were right it was D !!!! Thank you so much!!!!

OpenStudy (anonymous):

yw